OpenStudy (anonymous):

Help Please!!!!!!!!!!! Solve these equations using the addition property. x + y = 5 2x + y = 7 The lines intersect at ( __, __ )

OpenStudy (doc.brown):

I don't exactly know what addition property is. I did it like this, probably not what you're after...\[\left[\begin{matrix}1 & 1 & 5 \\ 2 & 1 & 7\end{matrix}\right]\] multiply row 1 by -2 and add it to row 2\[\left[\begin{matrix}1 & 1 & 5\\ 0 & -1 & -3\end{matrix}\right]\]multiply row 2 by -1\[\left[\begin{matrix}1 & 1 & 5\\ 0 & 1 & 3\end{matrix}\right]\]multiply row 2 by -1 and add to row 1\[\left[\begin{matrix}1 & 0 & 2\\ 0 & 1 & 3\end{matrix}\right]\]so x=2 y=3

OpenStudy (triciaal):

multiply the first eqn by -1 then add to the 2nd eqn (his will eliminate the y's ) solve for x substitute this value of x in either eqn to find y check your answer by substituting in either eqn

OpenStudy (triciaal):

-x-y = -5 (2x + y) + -(x + y) = 7 - 5 x = 2 x + y = 5 2 + y = 5 =>y = 3 check 2x + y = 7 2(2) + 3 =7 7=7 ok solution (2, 3)